Find an equation of the line containing the given pair of points (3,5) and (9,7)
Answer by Mathtut(3670)

find the slope
:
rise/run=m=7-5/9-3=2/6=1/3
:
point slope formula....using (3,5) as our point
:
y-5=(1/3)(x-3)
:
y-5=(1/3)x-1
:
y=(1/3)x+4